In a box containing blue and green beads, the number of blue beads are 2/3 to the number of green beads. What fraction of the total beards are blue.
Answer by greenestamps(13206)



Given that the number of blue beads is 2/3 the number of green beads, denote the number of green beads 3x and the number of blue beads 2x.

That makes the total number of beads 3x+2x=5x.

And then the fraction of the total number of beads that are blue is 2x/5x = 2/5.

ANSWER: 2/5 of the beads are blue.